P(A∪B)=61
or, 1−P(A∪B)=61
∴ P(A∪B)=1−61=65; P(A∩B)=41; and P(A)=41;
∴ P(A)=1−P(A)=1−41=43
We know, P(A∪B)=P(A)+P(B)−P(A∩B)
or, 65=43+P(B)−41
or, P(B)=65−21=31
Now, P(A).P(B)=43.31=41=P(A∩B)
i.e., events A and B are mutually independent.
Since the probability of A and B are different, so they are not equally likely events.
Therefore, (A) is the correct option.